OpenStudy (anonymous):

can someone please help me in algebra 1?

zepdrix (zepdrix):

Post a question c:

OpenStudy (anonymous):

Suppose y varies directly with x, and y = 25 when x = 140. What is the value of x when y = 36? (1 point) 140 165 201.6 176

zepdrix (zepdrix):

y varies directly with x,\[\large y=kx\]See how y and x have a direct relationship? They only differ by a constant, k. When x=140, y=25. We can use this information to find k.\[\large 25=k(140)\]To solve for k, divide both sides by 140.

zepdrix (zepdrix):

Which changes our equation to,\[\large y=kx \qquad \rightarrow \qquad y=\frac{25}{140}x\]When y=36, what is x?\[\large 36=\frac{25}{140}x\]To solve for x, multiply both sides by, \(\huge \frac{140}{25}\)